# Self-Managing DRAM: A Low-Cost Framework for Enabling Autonomous and Efficient in-DRAM Operations

### v6

@@ Abstract
We propose a new low-cost DRAM architecture, Self-Managing DRAM (SMD), that enables autonomous in-DRAM maintenance operations by transferring the responsibility for controlling maintenance operations from the memory controller to the SMD-chip. To enable autonomous maintenance operations, we tweak the DRAM interface with a single, simple modification, such that an SMD chip rejects memory controller accesses to DRAM regions (e.g., a subarray or a bank) under maintenance, while allowing memory accesses to other DRAM regions. Thus, SMD enables 1) implementing new in-DRAM maintenance mechanisms (or modifying existing ones) with no further changes in the DRAM interface, memory controller, or other system components, and 2) overlapping the latency of a maintenance operation in one DRAM region with the latency of accessing data in another.
We evaluate SMD and show that it 1) can be implemented without modifications to the physical DDRx interface with low latency (0.4% of row activation latency) and area (1.6% of a DRAM chip) overhead, 2) achieves 4.1% average speedup across 20 four-core memory-intensive workloads over a DDR4-based system/DRAM co-design technique that intelligently parallelizes maintenance operations with memory accesses, and 3) guarantees forward progress for rejected memory accesses.
@@ 1 Introduction
Advances in manufacturing technology enable increasingly smaller DRAM cell sizes, continuously reducing cost per bit of a DRAM chip [1, 2, 3, 4]. However, as a DRAM cell becomes smaller and the distance between adjacent cells shrink, ensuring reliable and efficient DRAM operation becomes an even more critical challenge [5, 6, 7, 8, 9, 10, 11, 12, 2, 13, 4, 14, 15]. A modern DRAM chip requires three types of maintenance operations (described in detail in section 5) for reliable and secure operation: 1) DRAM refresh [16, 17, 15, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], 2) RowHammer protection [44, 10, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 40, 56, 57, 58, 59, 60, 61, 62, 63], and 3) memory scrubbing.111Memory scrubbing is not always required in consumer systems but often used in cloud systems [64, 65, 66, 67, 68, 69, 70, 71]. New DRAM chip generations necessitate making existing maintenance operations more aggressive (e.g., lowering the refresh period [72, 73, 44]) and introducing new types of maintenance operations (e.g., targeted refresh [74, 75, 76] and DDR5 RFM [72] as RowHammer defenses).
@@
Two problems likely hinder effective and efficient maintenance mechanisms from being adopted in modern and future DRAM-based computing systems. First, it is difficult to modify existing maintenance mechanisms and introduce new maintenance operations because doing so often necessitates changes to the DRAM interface, which takes a long time (due to various issues related to standardization and agreement across many vendors). Second, it is challenging to keep the overhead of DRAM maintenance mechanisms low as DRAM reliability characteristics worsen and DRAM chips require more aggressive maintenance operations. We expand on the two problems in the next two paragraphs.
@@
Implementing new maintenance operations (or changing existing ones) often necessitates modifications in the DRAM interface, MC, and potentially other system components. Such modifications are only possible with a new DRAM standard, which takes a long time to develop, likely leading to slow progress in the adoption of new architectural techniques in DRAM chips. For example, there was a five-year gap between the DDR3 [77] and DDR4 [78] standards, and eight-year gap between the DDR4 [78] and DDR5 [72] standards. While developing a new standard, DRAM vendors need to push their proposal regarding the maintenance operations through the JEDEC committee (which is the case even if the gap between subsequent DRAM standards were to substantially reduce) such that the new standard includes the desired changes to enable new maintenance operations. Thus, a flexible DRAM interface that decouples improvements to DRAM maintenance operations from interface changes (and hence the timeline needed to enable such changes) would allow DRAM designers to quickly implement custom in-DRAM maintenance mechanisms and develop more reliable and secure DRAM chips.
@@
A maintenance operation triggers more frequently and takes longer with worsening DRAM reliability characteristics and increasing density. While a maintenance operation executes, DRAM is unavailable to serve memory requests (e.g., a periodic refresh operation in DDR4 prevents access to a DRAM rank for up to [78]). As a result, maintenance operations can significantly degrade system performance by delaying critical memory requests. Thus, enabling autonomous maintenance operations without delaying critical memory requests would improve system performance and energy efficiency.
@@
Our goal is to 1) ease and accelerate the process of implementing new in-DRAM maintenance operations and 2) enable more efficient maintenance operations. To this end, we propose Self-Managing DRAM (SMD), a new DRAM architecture that enables implementing new DRAM maintenance operations and modifying the existing ones with a single, simple interface change that eliminates the need for future changes in the DRAM interface, the MC, or other system components.
@@
SMD’s key idea is to allow a DRAM chip to autonomously and efficiently perform maintenance operations by preventing memory accesses only to a relatively small, under-maintenance DRAM region, i.e., a designated section in a DRAM chip (e.g., a DRAM subarray or a bank), while allowing memory accesses to other DRAM regions.222SMD has the key property that the MC does not even know which maintenance operations an SMD chip performs. This property of SMD could also be appealing to DRAM vendors because it would allow vendor-, and chip-generation-specific DRAM resilience characteristics (e.g., the degree and distribution of RowHammer vulnerability and retention failures) to remain undisclosed (i.e., not exposed to other DRAM vendors or system designers).
To prevent access to an under-maintenance DRAM region, an SMD chip rejects a row activation command (i.e., ACT) issued by the MC to the region. To ensure that a DRAM region cannot be locked for too long, SMD enforces a minimum delay (called ARI) between consecutive maintenance operations targeting the same region. We comprehensively study SMD and show that it ensures forward progress for memory requests (section 4.5).
While a DRAM region is under maintenance, the MC can access other regions. This way, a majority of memory accesses are not delayed by maintenance operations, as SMD overlaps the latency of a maintenance operation in a DRAM region with a memory accesses to another DRAM region.
@@
To enable practical adoption of SMD, we implement it with low-cost modifications in existing DRAM chips and MCs. First, to inform the MC that a row activation is rejected, SMD uses a single uni-directional pin on the DRAM chip (which already exists in DDR4/5, see section 7.4). Second, SMD implements a Lock Controller in a DRAM chip for managing regions under maintenance and preventing access to them. Third, SMD adds a new row address latch for every lock region in a DRAM chip to enable accessing one lock region while another is under maintenance (building on the basic design proposed in prior works [79, 80, 16]). In section 7.4, we show that these modifications have low DRAM chip area overhead (1.6% of a DRAM chip) and latency cost (only 0.4% additional row activation latency).
@@
We demonstrate the practicality and versatility of SMD by implementing in-DRAM maintenance mechanisms for DRAM refresh (section 5.1), RowHammer protection (section 5.2), and memory scrubbing (section 5.3). We rigorously evaluate SMD’s impact on system performance and energy efficiency using cycle-accurate memory system simulations (using Ramulator [81, 82]), executing a diverse set of 62 single-core and 60 four-core workloads. We compare SMD-based implementations of the evaluated maintenance mechanisms to their MC-based implementations. We make three key observations from our evaluation. First, an SMD chip that implements lock regions at subarray granularity (i.e., a maintenance mechanism locks a small designated section in a bank) and implements all three maintenance mechanisms (refresh, RowHammer protection, and scrubbing) 1) improves average performance across all memory-intensive four-core workloads by 8.6% compared to a baseline system with a DDR4 chip that supports bank-level refresh and a memory controller that intelligently schedules maintenance operations to avoid delaying main memory accesses (DARP [16]) and 2) outperforms a system in which the DDR4 chip can concurrently perform a maintenance operation in a subarray and a memory access in another subarray (maintenance-access parallelization) and the memory controller can intelligently exploit maintenance-access parallelization (e.g., similar to DSARP’s [16] refresh-access parallelization) by 4.1%.
Second, SMD (with maintenance-access parallelization) reduces DRAM energy consumption by 4.3% on average as it eliminates DRAM commands issued by the memory controller to perform maintenance operations and reduces total execution time. Third, a very low chip area overhead (0.001% of a DRAM chip) implementation of SMD that locks regions at bank granularity (i.e., a maintenance mechanism locks a DRAM bank) and performs only refresh operations for maintenance (similar to DDR4) induces 4.5% average slowdown due to rejected row activations across all memory-intensive four-core workloads compared to the same baseline system (DARP [16]). We conclude that SMD practically provides better performance and energy benefits, without having to modify the DRAM interface (e.g., by adding new timing parameters to the standard) for every new type of maintenance mechanism, than system-DRAM co-design techniques that enable refresh-access parallelization (e.g., [16, 80]).
@@
We expect and hope that SMD will inspire researchers to develop new DRAM maintenance mechanisms that more efficiently tackle the reliability and security problems of DRAM and to enable practical adoption of innovative ideas by providing freedom to both DRAM and MC designers. We will publicly release all SMD source code and data so that others can replicate and build on the work.
@@
We propose SMD, a new DRAM chip design and interface to enable autonomous and efficient in-DRAM maintenance operations. We implement SMD with small changes to modern DRAM chips and MCs.
We use SMD to implement efficient DRAM maintenance mechanisms for three use cases: DRAM refresh, RowHammer protection, and memory scrubbing.
We rigorously evaluate the performance and energy of SMD-based maintenance mechanisms. SMD provides large performance and energy benefits while also improving reliability across a variety of systems and workloads.
@@ 2.1 DRAM Organization
A computing system has one or more DRAM channels, where each channel has an independently operating I/O bus. As Fig. 1 illustrates, a Memory Controller (MC) interfaces with one or multiple DRAM ranks via the channel’s I/O bus. A DRAM rank consists of a group of DRAM chips that operate in lockstep. Because the I/O bus is shared across ranks, accesses to different ranks happen in serial manner. A DRAM chip is divided into multiple DRAM banks, each of which is further divided into multiple two dimensional DRAM cell arrays, called subarrays. Within each subarray, DRAM cells are organized as rows and columns. A DRAM row consists of hundreds of DRAM cells which connect to the same wordline. A DRAM cell connects to a sense amplifier via a bitline, and all sense amplifiers in the subarray form a row buffer. Within each DRAM cell, the data is stored as electrical charge on the capacitor and accessed through an access transistor.
@@ 2.2 Accessing DRAM
A row must be activated (i.e., opened), to load the row’s data into the row buffer. A row activation is performed by enabling the wordline corresponding to the row address of the ACT command. To enable the corresponding wordline, the global row decoder first partially decodes the row address, and the local row decoder selects a wordline based on the partially decoded address provided by the global row decoder. Enabling the wordline copies the row’s data to the row buffer.
Data in the row buffer is accessible by a RD/WR. The MC follows an ACT with one or multiple RD/WR commands at least after the activation latency (tRCD). The MC issues a PRE command to close a row, after which it can activate a new row from the same bank. The time interval for issuing a PRE after an ACT must be at least restoration latency (tRAS). The MC follows a PRE with an ACT to activate a new row at least after the precharge latency (tRP).
3 Motivation
In current DRAM chips, the MC is in charge of managing DRAM maintenance operations such as periodic refresh, RH protection, and memory scrubbing. When DRAM vendors modify a DRAM maintenance mechanism, the changes often need to be reflected to the MC design as well as the DRAM interface, which makes such modifications very difficult. As a result, implementing new or modifying existing maintenance operations, no matter how fast such implementations or modifications could be developed, can only be realized after a multi-year effort by multiple parties that are part of the JEDEC committee. A prime example to support our argument is the most recent DDR5 standard [72], which took almost a decade to develop after the initial release of DDR4. Even though it might not have taken DRAM designers a decade to develop new maintenance mechanisms, the mechanisms are implemented only after the new standard is released. DDR5 also introduces changes to key issues we study in this paper: DRAM refresh, RH protection, and memory scrubbing. We discuss these changes as motivating examples to show the shortcomings of the status quo in DRAM.

### v7

# Self-Managing DRAM: A Low-Cost Framework for Enabling Autonomous and Efficient DRAM Maintenance Operations

@@ Abstract
We propose a new low-cost DRAM architecture, Self-Managing DRAM (SMD), that enables autonomous in-DRAM maintenance operations by transferring the responsibility for controlling maintenance operations from the memory controller to the SMD chip. To enable autonomous maintenance operations, we make a single, simple modification to the DRAM interface, such that an SMD chip rejects memory controller accesses to DRAM regions (e.g., a subarray or a bank) under maintenance, while allowing memory accesses to other DRAM regions. Thus, SMD enables 1) implementing new in-DRAM maintenance mechanisms (or modifying existing ones) with no further changes in the DRAM interface, memory controller, or other system components, and 2) overlapping the latency of a maintenance operation in one DRAM region with the latency of accessing data in another.
We evaluate SMD and show that it 1) can be implemented without adding new pins to the DDRx interface with low latency (0.4% of row activation latency) and area (1.1% of a DRAM chip) overhead, 2) achieves 4.1% average speedup across 20 four-core memory-intensive workloads over a DDR4-based system/DRAM co-design technique that intelligently parallelizes maintenance operations with memory accesses, and 3) guarantees forward progress for rejected memory accesses. We believe and hope SMD can enable innovations in DRAM architecture to rapidly come to fruition. We open source all SMD source code and data at https://github.com/CMU-SAFARI/SelfManagingDRAM.
IEEEexample:BSTcontrol
@@ 1 Introduction
Advances in manufacturing technology enable increasingly smaller DRAM cell sizes, continuously reducing cost per bit of a DRAM chip [1, 2, 3, 4, 5, 6]. However, as a DRAM cell becomes smaller and the distance between adjacent cells shrinks, ensuring reliable and efficient DRAM operation becomes an even more critical challenge [7, 8, 9, 10, 11, 12, 13, 14, 2, 15, 5, 16, 17, 18]. A modern DRAM chip requires three major types of maintenance operations (described in detail in section 5) for reliable and secure operation: 1) DRAM refresh [19, 20, 17, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], 2) RowHammer protection [47, 48, 49, 50, 12, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 18, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122], and 3) memory scrubbing [123, 124, 125, 126, 127, 128, 129, 130, 131, 17, 132, 133, 134, 135].111Memory scrubbing is not always required in consumer systems but often used in cloud systems. New DRAM chip generations necessitate making existing maintenance operations more aggressive (e.g., lowering the refresh period [119, 136, 137]) and introducing new types of maintenance operations (e.g., targeted refresh [66, 64, 138], DDR5 RFM [119], and PRAC [119] as RowHammer defenses).222 A very recent update to the DDR5 standard [119] introduces PRAC, which is an on-DRAM-die read disturbance mitigation mechanism. PRAC requires more changes to the DRAM interface and continues to use RFM. Note that PRAC is concurrent with this work, as the initial version of this paper [139] was placed on arXiv on 27 July 2022 and initial submission to the MICRO 2022 conference was made on 22 April 2022.
@@
Two problems likely hinder the adoption of effective and efficient maintenance mechanisms in modern and future DRAM-based computing systems. First, it is difficult to modify existing maintenance mechanisms and introduce new maintenance operations because doing so often necessitates changes to the DRAM interface, which takes a long time (due to various issues related to standardization and agreement across many vendors with conflicting interests [4, 6]). Second, it is challenging to keep the overhead of DRAM maintenance mechanisms low as DRAM reliability characteristics worsen and DRAM chips require more aggressive maintenance operations. We expand on the two problems in the next two paragraphs.
@@
Implementing new maintenance operations (or changing existing ones) often necessitates modifications in the DRAM interface, memory controller (MC), and potentially other system components. Such modifications are only possible with a new DRAM standard, which takes a long time to develop, likely leading to slow progress in the adoption of new architectural techniques in DRAM chips. For example, there was a five-year gap between the DDR3 [140] and DDR4 [141] standards, and an eight-year gap between the DDR4 [141] and DDR5 [119] standards. While developing a new standard, DRAM vendors need to push their proposal regarding the maintenance operations through the JEDEC committee (which is the case even if the gap between subsequent DRAM standards were to substantially reduce) such that the new standard includes the desired changes to enable new maintenance operations. Thus, a flexible DRAM interface that decouples improvements to DRAM maintenance operations from interface changes (and hence the timeline needed to enable such changes) would allow DRAM designers to quickly implement custom in-DRAM maintenance mechanisms and develop more robust, reliable, secure, safe, and more efficient DRAM chips.
@@
A maintenance operation triggers more frequently and takes longer with worsening DRAM reliability characteristics, increasing density, or increasing operating temperature with tighter DRAM-system integration (e.g., high-bandwidth memory DRAM temperature can reach [142]). While a maintenance operation executes, DRAM is unavailable to serve memory requests (e.g., a periodic refresh operation in DDR4 prevents access to a DRAM rank for [141] and in DDR5 prevents access to a DRAM bank in all bank groups for [119]). As a result, maintenance operations can significantly degrade system performance by delaying critical memory requests. Enabling autonomous maintenance operations without delaying critical memory requests would improve system performance, energy efficiency, and robustness.
Enabling autonomous operations with a flexible interface would provide two main benefits. First, it would enable DRAM manufacturers with breathing room to perform architectural optimizations that can greatly benefit from their DRAM-internal proprietary knowledge without having to expose such knowledge to off-chip components and third-party manufacturers, which is impractical [4, 6]. Second, doing so divides the work nicely between the memory controller and the DRAM chip. DRAM manufacturers can focus on their critical challenges of designing better DRAM chips and memory controller manufacturers can focus on the critical tasks for memory controller without being burdened by other tasks that are not really easy to optimize at the memory controller level without proprietary DRAM-internal knowledge.
@@
Our goal is to 1) ease and accelerate the process of implementing new in-DRAM maintenance operations and 2) enable more efficient maintenance operations. To this end, we propose Self-Managing DRAM (SMD), a new framework that enables implementing new in-DRAM maintenance operations and modifying the existing ones with a single, simple interface change that eliminates the need for future changes in the DRAM interface, the MC, or other system components.333The current commodity DRAM interface (e.g., DDRx [140, 141, 119], HBMx [143, 144]) is completely processor centric (i.e., the memory controller on the processor chip dictates everything to the DRAM chips, giving no breathing room for DRAM to do much on its own). SMD, by slightly modifying the interface, enables the interface to move towards and more memory-centric approach where memory can do things autonomously. This is in line with the general memory-centric computing paradigm [145, 146, 147] where memory can also do computation.
@@
SMD’s key idea is to allow a DRAM chip to autonomously and efficiently perform maintenance operations by preventing memory accesses only to a relatively small, under-maintenance DRAM region, i.e., a designated section in a DRAM chip (e.g., a DRAM subarray or a bank), while allowing memory accesses to other DRAM regions.444SMD has the key property that the MC does not even know which maintenance operations an SMD chip performs. This property of SMD could also be appealing to DRAM vendors [4, 6] because it would allow vendor-, and chip-generation-specific DRAM resilience characteristics (e.g., the degree and distribution of RowHammer vulnerability and retention failures) to remain undisclosed (i.e., not exposed to other DRAM vendors or system designers). To prevent access to an under-maintenance DRAM region, an SMD chip rejects a row activation command (i.e., ACT) issued by the MC to the region. To ensure that a DRAM region cannot be locked for too long, SMD enforces a minimum delay between consecutive maintenance operations targeting the same region. We comprehensively study SMD and show that it ensures forward progress for memory requests (section 4.4). While a DRAM region is under maintenance, the MC can access other regions. This way, a majority of memory accesses are not delayed by maintenance operations as SMD overlaps the latency of a maintenance operation in a DRAM region with memory accesses to other DRAM regions.
@@
To enable practical adoption of SMD, we propose to implement it with low-cost modifications to existing DRAM chips and MCs. First, to inform the MC that a row activation is rejected, SMD uses a single uni-directional pin on the DRAM chip (which already exists in DDR4/5, see section 6). Second, SMD implements a small (e.g., 16 bits per DRAM bank) memory structure called the Lock Controller in a DRAM chip for managing regions under maintenance and preventing access to them. Third, SMD adds a new row address latch for every lock region in a DRAM chip to enable accessing one lock region while another is under maintenance (building on the basic design proposed in prior works [148, 149, 19]). In section 6, we show that these modifications have low DRAM chip area (1.1% of a DRAM chip) and negligible latency overhead (only 0.4% increase in row activation latency).
@@
We demonstrate the usefulness and versatility of SMD by implementing three in-DRAM maintenance mechanisms for DRAM refresh (section 5.1), RowHammer protection (section 5.2), and memory scrubbing (section 5.3). We rigorously evaluate SMD’s impact on system performance and energy efficiency using cycle-accurate memory system simulations (using Ramulator [150, 151]), executing a diverse set of 62 single-core and 60 four-core workloads. We compare SMD-based implementations of the evaluated maintenance mechanisms to their MC-based implementations. We make two key observations from our evaluation. First, an SMD chip that implements lock regions at subarray granularity (i.e., a maintenance mechanism locks a small designated section in a bank) and implements all three maintenance mechanisms (refresh, RowHammer protection, and scrubbing) 1) improves average performance across all memory-intensive four-core workloads by 8.6% compared to a baseline system with a DDR4 chip that supports bank-level refresh and a memory controller that intelligently schedules maintenance operations to avoid delaying main memory accesses (DARP [19]) and 2) outperforms an optimized state-of-the-art system in which the DDR4 chip can concurrently perform a maintenance operation in a subarray and a memory access in another subarray (maintenance-access parallelization) and the memory controller can intelligently exploit maintenance-access parallelization (e.g., similar to DSARP’s [19] refresh-access parallelization) by 4.1%. SMD provides 88.3% of the speedup provided by an oracle (No-Refresh) that completely mitigates the overhead of maintenance operations, on average across all evaluated high memory intensity four core workloads. Second, SMD reduces DRAM energy consumption by 4.3% on average as it eliminates DRAM commands issued by the memory controller to perform maintenance operations and reduces total execution time. We conclude that SMD practically provides better performance and energy benefits, without having to modify the DRAM interface (e.g., by adding new timing parameters to the standard) for every new type of maintenance mechanism, than prior techniques that enable refresh-access parallelization (e.g., [19, 149]).
@@
We expect and hope that SMD will inspire researchers to develop new DRAM maintenance mechanisms that more efficiently tackle the scaling, robustness, and efficiency problems of DRAM and to enable practical adoption of innovative ideas by providing freedom to both DRAM and MC designers. We release all SMD source code and data at https://github.com/CMU-SAFARI/SelfManagingDRAM so that others can replicate our work and build on it.
@@
We propose SMD, a new framework to enable autonomous and efficient in-DRAM maintenance operations with small changes to modern DRAM chips and MCs.
We use SMD to implement efficient DRAM maintenance mechanisms for three use cases: DRAM refresh, RowHammer protection, and memory scrubbing. We also show that SMD can seamlessly enable optimized maintenance mechanisms, including variable-rate refresh and probabilistic RowHammer protection techniques.
We rigorously evaluate the performance and energy of SMD-based maintenance mechanisms. SMD provides large performance and energy benefits while also improving system robustness across a variety of workloads.
@@ 2.1 DRAM Organization
A computing system has one or more DRAM channels, where each channel has an independently operating I/O bus. As Fig. 1 illustrates, a Memory Controller (MC) interfaces with one or multiple DRAM ranks via the channel’s I/O bus. A DRAM rank consists of a group of DRAM chips that operate in lockstep. Because the I/O bus is shared across ranks, memory accesses to different ranks happen in serial manner. A DRAM chip is divided into multiple DRAM banks, each of which is further divided into multiple two dimensional DRAM cell arrays, called subarrays. Within each subarray, DRAM cells are organized as rows and columns. A DRAM row consists of DRAM cells that are connected to the same wordline. A DRAM cell connects to a sense amplifier via an access transistor and a bitline, and all sense amplifiers in the subarray form a row buffer. Within each DRAM cell, the data is stored as electrical charge on the capacitor.
@@ 2.2 Accessing DRAM
Reading or updating the value of a DRAM cell requires activating (i.e., opening) a DRAM row. To open a row, the memory controller issues a DRAM activate (ACT) command with a row address. The global row decoder and the local row decoder select a wordline based on the row address provided by the memory controller. Enabling the wordline copies the row’s data to the row buffer.
The memory controller can access data in the row buffer using DRAM read and write commands (RD/WR). The delay between an ACT and the following RD or WR must be larger than activation latency (tRCD). The memory controller can issue a PRE command to close the open row, after which it can activate a new row from the same bank. The memory controller must wait for at least the restoration latency (tRAS) after issuing an ACT before issuing a PRE. The memory controller follows a PRE with an ACT to activate a new row at least after precharge latency (tRP).
3 Motivation & Goal
In current DRAM chips, the MC is in charge of managing DRAM maintenance operations such as periodic refresh, RowHammer protection, and memory scrubbing. When DRAM vendors modify a DRAM maintenance mechanism to improve robustness, performance, or efficiency, often the DRAM interface needs to be modified as well, which makes such improvements very difficult to achieve. As a result, implementing new or modifying existing maintenance operations, no matter how fast such implementations or modifications could be developed, can only be realized after a multi-year effort by multiple parties that are part of the JEDEC committee [6, 4, 152]. A prime example to support our argument is the most recent DDR5 standard [119], which took at least 8 years to develop after the initial release of DDR4. Even though it might not have taken DRAM designers so long to develop new maintenance mechanisms, the DRAM chips implementing these mechanisms are released only after the new standard is released. DDR5 also introduces changes to key issues we study in this paper: DRAM refresh, RH protection, and memory scrubbing. We discuss these changes as motivating examples to show the shortcomings of the status quo in DRAM.